### v1

@@ Acknowledgements.
Software: JHelioviewer  [^3] https://www.jhelioviewer.org/, has been used as a visualisation tool for solar data from different instruments (Müller et al., 2017). This research has made use of PyThea v0.7.4, an open-source and free Python package to reconstruct the 3D structure of CMEs and shock waves (Zenodo: https://doi.org/10.5281/zenodo.5713659). We also used GCS in Python v0.2.2 (Zenodo: https://zenodo.org/badge/latestdoi/297350666), which is a python 3 implementation of the Graduated Cylindrical Shell model based on the existing IDL implementation in SolarSoft. Other python(v3.9) libraries used for this work are: astropy (Astropy Collaboration et al. (2018)), matplotlib (Hunter (2007)), numpy (van der Walt et al. (2011)), pandas (Reback et al. (2020)), scipy (Virtanen et al. (2020)), SunPy (SunPy Community et al. (2020)). The SolarSoft  [^4] https://soho.nascom.nasa.gov/solarsoft/ package in IDL(v8.6.0) has been used in this paper. Also the WISPR package on Solarsoft has been widely used.

### v2

@@ Acknowledgements.
Software: JHelioviewer  [^4] https://www.jhelioviewer.org/, has been used as a visualization tool for solar data from different instruments (Müller et al. 2017). This research has made use of PyThea v0.7.4, an open-source and free Python package to reconstruct the 3D structure of CMEs and shock waves (Zenodo: https://doi.org/10.5281/zenodo.5713659). We also used GCS in Python v0.2.2 (Zenodo: https://zenodo.org/badge/latestdoi/297350666), which is a python 3 implementation of the Graduated Cylindrical Shell model based on the existing IDL implementation in SolarSoft. Other python(v3.9) libraries used for this work are: astropy (Astropy Collaboration et al. (2018)), matplotlib (Hunter (2007)), numpy (van der Walt et al. (2011)), pandas (Reback et al. (2020)), scipy (Virtanen et al. (2020)), SunPy (SunPy Community et al. (2020)). The SolarSoft  [^5] https://soho.nascom.nasa.gov/solarsoft/ package in IDL(v8.6.0) has been used in this paper. Also the WISPR package on Solarsoft has been widely used.